Find the standard form of the equation of the line with slope 3/5 and y-intercept -6.
Found 2 solutions by Mathtut, stanbon:
Answer by Mathtut(3670)

y=(3/5)x-6
:
multiply each term by 5
:
5y=3x-30
:
3x-5y=30 is standard form

Answer by stanbon(75887)

Find the standard form of the equation of the line with slope 3/5 and y-intercept -6.
------------
y = (3/5)x - 6
5y = 3x - 30
3x - 5y = 30
===============
Cheers,
Stan H.